LUCIFER Commissioning and AGw Commissioning for 14-15 September 2008

Observers: J. Hill, N. Ageorges, W. Seifert, and numerous other LUCIFERians (Peter, Kai, Marcus, Volker, Andres) on mtn
More Observers: D. Miller in Tucson and A. Rakich (second half)
Software Support: T. Liebold in Tucson (first half)
Telescope Operators: D. Gonzalez Huerta and W. Wack

Summary

Lost first half of night to rain, clouds, fog. Used 1.5 hours to debug LUCIFER motor controller problems. Then used 1.5 hours on sky to confirm that GCS does the correct wavefront sensing in the off-Y-axis positions. Back to zenith at 3AM to work on LUCIFER, used 45 min to diagnose and remove failed power supply. Used remained of night to collect data on M2 alignment. Struggled with some minor problems.

Note that LUCIFER will need telescope at Zenith from 3PM to 6PM Monday to reinstall and test their repaired power supply.

Nighttime Activities

Rain and Fog

01:00 Heavy rain and thunderstorm blowing over us from the Northeast.

02:00 Merged the -1.533 degrees of LEFTIAA into LEFTZEROPOINT, to correct the angles of combined (probe plus telescope) offsets as measured by LUCIFER last night. Now LEFTIAA=0.0 and LEFTZERPOINT=-187.433. PCS has been restarted.

02:08 Restarted irs and IIF in coordination to try to get rid of irs hanging after just a few GetMultiParams from LUCIFER.

03:00 Rain has stopped and sky has cleared. Now we are waiting for things to dry off.

Norm and Wayne and John were doing various experiments on irs and IIF. One time, irs got hung when we changed the TCS authorization to LBC, but in a repeat experiment it worked fine switching back and forth from LUCIFER to LBC.

04:30 Torsten has installed a new version of GCS on the mountain. This one should be able to make correct active optics corrections from any probe position.

05:00 Building has mostly dried off, but now we have fog blowing over the mountain from the North.

05:30 Walter is doing LUCIFER flexure tests with dome closed. The present mystery is why the flexure seen here is not repeatable as was the flexure measured in Heidelberg on the test stand. Apparently one of the pinhole masks is a bit loose.

06:30 John is rough pumping the Belljar to store it under vacuum.

Open for Observing

07:15 Fog has cleared and humidity is down to 88%, so we are going to open. T=6 degC, D=4 degC.

07:30 Sitting at zenith focussing on stars drifting by while Peter debugs a LUCIFER issue. Doug is testing the M2 anemometer.

08:30 LUCIFER still having motor controller problems, so we take an hour for GCS and alignment tests.

GCS WFS Testing

08:35 pointto M5_262 but there was confusion about sidereal time.

08:43 pointto BS9107, found at 75,217 with default pointing model; centered with IE=-80 CA=-57. (apparently the wrong star)

08:48 restarted hung IRS.

08:53 pointto BS9107 ACTIVE on-axis

08:57 Stop GCS and restart so we get the new one. (but guide rate is still 0.4 Hz) PCS guiding was momentarily blocked.

08:58 pointto BS9107 ACTIVE on-axis up through wfsc 817

09:04 pointto BS9107 off-axis gs=91, x=0 y=499.3 but no star even in 1 sec acq exposure.

09:10 pointto M5_1133 on-axis TRACK now really centered at: IE=-40 CA=-55

09:15 back to BS9107 TRACK - star is there

09:18 pointto BS9107 off-axis gs=91 (too faint for 100 ms, increase to 1000 ms), collimating off-axis first image was wfsc 820

09:20 GCS apparently stopped WFSing when it couldn't see the star. That's good, but we couldn't get it to resume.

09:26 pointto BS9107 off-axis gs=35 ACTIVE - WFS didn't start up. Was flag not reset?

09:27 stop and restart GCS

09:26 pointto BS9107 off-axis gs=35 ACTIVE PA=315 (on-Y-axis) wfsc through 825 (has the quadrifoil effect in WFS spots

09:32 pointto BS9107 off-axis gs=35 PA=265 probe x=114 Y=518

Noticing some 2 arcsec jumps in guiding image.

Problem with WFS loop stopping. Had to say stopWFS, and then triggerWFS

09:43 Wayne presses startWFS button on GCSGUI (This restarted loop.) wfsc images through 0831

09:45 pointto BS9107 on-axis TRACK

Apply -25 arcsec Tilt CA=-30

09:51 We crashed GCS, so restart. The small corrections applied here at the on-axis star mean that we correctly converged for the off-Y-axis star.

02:53:32 ACTIVE mode took 2 WFS image and then stopped taking more images.

09:55 Wayne does stopWFS and startWFS buttons, but the hung WFS loop doesn't stop

09:59 Put tilt and CA back to -55. Didn't get enough time

LUCIFER Work

10:00 Close dome for zenith work on LUCIFER

Had to kill GCS as it wouldn't stop guiding as we went to horizon to close dome.

A 36-volt power supply had failed on LUCIFER, so this is a nice explanation why their motor controllers wouldn't work. They have dismounted the power supply, and taken it downstairs for repair.

M2 Alignment Measurements

10:50 Reopen the dome to continue with secondary alignment measurements.

Roof wind measurement reads 15 m/sec with dome closed, but only 8 m/sec with dome open.

10:58 pointto BS9107 on-axis in ACTIVE mode. (RPC error because Wayne had killed GCS previously.)

11:00 pointto BS9107 on-axis in ACTIVE mode. setGuideRate 0.1 since star is near

Sep 15 04:01:57 lbtmu105 LBT_IIF: [I] PCS set new target to RA: 9.252426697134940e-01 Dec: 6.517835128836602e-01 complete
Sep 15 04:01:57 lbtmu105 LBT_IIF: [I] Invalid buffer: RPCERROR

11:05 Stop PCS and MCS and restart, although we can't really tell who made the RPC error.

11:07 Stop IIF and irs.

11:09 Send pointto again.

Sep 15 04:09:34 lbtmu105 LBT_IIF: [I] PCS set new target to RA: 9.252426697134940e-01 Dec: 6.517835128836602e-01 complete
Sep 15 04:09:34 lbtmu105 LBT_IIF: [I] Invalid buffer: RPCERROR

11:10 Stop and restart PSFL (PSFL was the guilty partly in this case)

11:11 Send pointto again. Zglobal=-0.1 IE=-55 CA=-30 (was guide probe in wrong position?)

11:21 IE=-35 CA=-35

11:22 pointto BS9107 on-axis ACTIVE starting with lots of coma at wfsc 835 setGuideRate 0.1 continue through wfsc 844

11:30 pointto BS9107 off-axis gs=69 GUIDE

11:34 pointto BS9107 on-axis GUIDE

11:35 Send -25 tilt on secondary and -25 in CA. Lost star

11:39 pointto BS9107 TRACK reset IE=-30 CA = -60

11:46 tilt -5 moved up CA delta +5

11:47 tilt -25, CA -35

11:49 GCS froze, Wayne killed GCSL

11:50 pointto BS9107, GUIDE collimate 2-7

11:53 gstar=69 wfsc 7-10 all failed after acquisition due to filename in IDL. Fixed

12:07 on-axis, collimate 11-14

12:13 off-axis, gstar 69, wfsc 15-16 (had removed field corrections, will fix later)

12:17 on-axis, collimate 17

12:20 off-axis, gstar 29, wfsc 18-19

12:23 on-axis, collimate 20-21

12:27 off-axis, gstar 87 22-23

12:33 Daylight! Done. Close dome.

Important Note about IE and CA jumps that we saw above.

When I was recording the collimation offsets at the end of the night, I found M1 pointing offsets that should not have been there. I think this must be the explanation for the mysterious jump in IE that we saw between 11:11 and 11:39.

M1 Pointing Offsets X=0 Y=-1.396 Z=0 RX=-30 RY=0 (not expected)

M2 Pointing Offsets X=+1.277 Y=0 Z=0 RX=0 RY=+245.5 (-25 arcsec Tilt)

M2 Global Offsets X=-7.287 Y=-3.266 Z=0 RX=222.3 RY=-63.7

M1 Global Offsets X=-0.180 Y=+1.100 Z=-0.1 RX=+13.7 RY=+13.3

M1 Active Corrections X=-1.264 Y=-0.437 Z=-0.309 RX=-4.7 RY=+13.6

Updating the various tables:

The new value of CA including the -25 arcsec of M2 Tilt should be: CA=-35
The value of IE without the above stray M1 Pointing Offsets should be: IE=-60
These have been placed in LUCIFERSX.20080911.ptmod and irtc.20080914.ptmod with PCSInstrument.conf updated appropriately.

The new M2 global offsets with the -25 arcsec Tilt included X=-6.010 Y=-3.266 Z=0 RX=+222.3 RY=+181.8
These have been placed in SXSMIRTCFCollimation.20080914.dat and SXSMLUCIFER.20080914.dat.

The new M1 global offsets with the last Active Corrections included: X=-1.444 Y=+0.663 Z=-0.409 RX=+9.0 RY=+26.9
These have been placed in SXPMIRTCFCollimation.20080914.dat and SXPMLUCIFER.20080914.dat.

-- JohnHill - 15 Sep 2008
Topic revision: r9 - 26 Sep 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback